OpenStudy (anonymous):

Complete the conversion, round to the nearest hundreths. 75m/h ≈ cm/h

OpenStudy (anonymous):

m = 100cm 1 = m/100cm = 100cm/m Set it up the way I set up the last problem.

OpenStudy (anonymous):

could u explain it one more time plz

OpenStudy (anonymous):

Anything times itself is 1, so if we multiply the thing we wanna convert by m/100cm or 100cm/m it will stay the same because they both = 1. If we choose the second one, 100cm/m, then the meters will cancel out because it's on the top and the bottom. So, 75m/h * 100cm/m = 75*100cm/h 7500cm/h

OpenStudy (anonymous):

thank u